OpenStudy (anonymous):

A coin is tossed and a die is rolled. What is the probability that the coin shows tails and the die shows 3?

OpenStudy (anonymous):

Coin: 1/2 Die: 3 = 1 choice out of 6 = 1/6 Multiply the two because of the AND : 1/2* 1/6 = 1/12